OpenStudy (anonymous):

Given the following function, y = f(x), find the indicated values: f(0), f(-1), f(-3)

OpenStudy (ranga):

The problem is incomplete. What is the function f(x)?

OpenStudy (anonymous):

just substitute that value for every x in the function. So if your funtion is 4x-3, then for f(-1) just put it as: 4(x)-3 = 4(-1)-3 = -4-3 = -7 f(-3) = 4(x)-3 = 4(-3)-3 = -12-3 = -15 :)

OpenStudy (anonymous):

here is the graph

OpenStudy (anonymous):

sorry the first graph is to a different problem, the 2nd graph is right

OpenStudy (anonymous):

ok So you have to match up where the x coordinates are according to the y coordinates, so when f(0), that means your x=0, so when x=0, your y value is at y=-5

OpenStudy (anonymous):

Unless they want you to make an equation out of the graph and then solve

OpenStudy (anonymous):

When it asks for f(0) would it be -5 or (-3,1) Im confused

OpenStudy (anonymous):

oh ok so it would be the -5

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

so for f(-1) it would be x=-7

OpenStudy (anonymous):

yes :)

OpenStudy (ranga):

Find f(0) means what is the y value when x = 0

OpenStudy (anonymous):

and f(-3) would be x=0

OpenStudy (anonymous):

but remember its y you are looking for not x

OpenStudy (anonymous):

So your typing f(-3) is x=0, but it should be y=0

OpenStudy (ranga):

f(-3) means what is the y value when x = -3?

OpenStudy (anonymous):

but actually when f(-3), then y=1

OpenStudy (anonymous):

ok great I think I figured it...thank you guys so much

OpenStudy (anonymous):

just remember that f(x)= means the same thing as y=

OpenStudy (anonymous):

yea y=1 when f(-3)

OpenStudy (anonymous):

Exactly !

OpenStudy (anonymous):

cool great

OpenStudy (anonymous):

that helped a lot

OpenStudy (ranga):

What do you get for f(-1)?

OpenStudy (anonymous):

y=-7

OpenStudy (ranga):

You got it!

OpenStudy (anonymous):

heehaw, lol

OpenStudy (anonymous):

Yay good job :)

OpenStudy (anonymous):

thanks again for the help, you all have a good one

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

no problemo!